OpenStudy (anonymous):

-n/15 - 4 great than or equal to -4

OpenStudy (texaschic101):

-n/15 - 4 >= -4 -n/15 >= -4 + 4 -n/15 >= 0 (multiply the equation by 15) -n >= 0 n >= 0 I am not sure about this one.....I must be doing something wrong. You need to get somebody else over here to see if they can help you...sorry

OpenStudy (anonymous):

makes sense i try it out and see if it fits. thanks

OpenStudy (mertsj):

\[\frac{-n}{15}-4 \ge -4\] \[\frac{-n}{15}\ge 0\] \[-n \ge 0\] \[n \le 0\]

OpenStudy (anonymous):

Thanks, thats what I got but was unsure.

OpenStudy (mertsj):

yw